# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ I Introduction
To protect QKD systems from eavesdropping attacks, researchers have made significant contributions. For example, measurement-device-independent (MDI) QKD was proposed, removing the security assumptions associated with the measurement and ensuring that the security of the QKD system is no longer threatened by potential vulnerabilities in the measurement unit [26]. Additionally, the decoy-state protocol was proposed to enhance the security performance and efficiency of a QKD system with phase-randomized weak coherent sources [27, 28]. While enhancing the security of QKD systems, efforts have also been made to improve the key rate of the system by raising the repetition rate of key distribution. The exploration of high-speed QKD systems has become a prominent area of research.
@@ III Security proof with intensity correlations
As mentioned at the end of Sec. II, our goal is to quantitatively provide the lower bound of single-photon yield $Y_{11,ab}^{k,\text{act}}$ and the upper bound of single-photon error probability $H_{11,ab}^{k,\text{act}}$. Based on the convexity of $G_{-}$ and concavity of $G_{+}$, Eq. 8 can be further scaled as
@@
Substituting three intensity settings $\mu,\nu,\omega$ into Eq. 13, the lower bound of $Y_{11}^{\text{ref}}$ and the upper bound of $H_{11}^{\text{ref}}$ can be solved by Gaussian elimination, which are expressed as [38]
@@ IV Key rate simulation
In the simulation, $Q_{ab}$ and $E_{ab}$ depend on Alice and Bob’s intensity settings $a$ and $b$, the total transmittance of the channel and detector $\eta$, the probability of background counts $Y_{0}$, and misalignment errors $e_{d}$. Appendix B provides the calculations of $Q_{ab}$ and $E_{ab}$ using these parameters. Consistent with Refs. [34, 35], since we are considering the asymptotic regime, the key rate does not depend on the probabilities of decoy settings and basis choices. Therefore, we setting $p_{\mu}\approx 1$ and $q_{Z}\approx 1$ to maximizes key rate similar to Ref. [35]. The other simulation parameters are utilized from Ref. [40] as follows. The bidirectional error correction efficiency to $f=1.16$, the probability of background counts $Y_{0}=4\times 10^{-8}$, and misalignment errors $e_{d}=0.0108$. For the total transmittance of the channel and detector, $\eta=\eta_{d}10^{-\alpha L/10}$, where $\eta_{d}$ is the detection efficiency of the detector, $\alpha$ is the attenuation coefficient of the channel, and $L$ is the distance from Alice (Bob) to Charles. In the simulation, we set $\eta_{d}=0.53$ and $\alpha=0.2$ dB/km.
@@ V Application
It is worth noting that, according to the security proof of decoy-state protocol [27, 28], the key rate obtained using the signal state, weak decoy state, and vacuum state can approximate the key rate obtained with an infinite number of decoy states. Therefore, to simplify calculations while maintaining accuracy, we choose intensities of signal state $S$, weak decoy state $D_{1}$, and vacuum state $V$ in our simulation. Next, taking the signal state $S$ as an example, we analyze the relative deviation $\delta$ under intensity correlations.

### v2

@@ I Introduction
To protect QKD systems from eavesdropping attacks, researchers have made significant contributions. For example, measurement-device-independent (MDI) QKD was proposed, removing the security assumptions associated with the measurement and ensuring that the security of the QKD system is no longer threatened by potential vulnerabilities in the measurement unit [26]. Additionally, the decoy-state protocol was proposed to enhance the security performance and efficiency of a QKD system with phase-randomized weak coherent sources [27, 28, 29]. While enhancing the security of QKD systems, efforts have also been made to improve the key rate of the system by raising the repetition rate of key distribution. The exploration of high-speed QKD systems has become a prominent area of research.
@@ III Security proof with intensity correlations
As mentioned at the end of Sec. II, our goal is to quantitatively provide the lower bound of single-photon yield $Y_{11,ab}^{k,\text{act}}$ and the upper bound of single-photon error probability $H_{11,ab}^{k,\text{act}}$. Based on the convexity of $G_{-}$ and concavity of $G_{+}$q. 8 can be further scaled as
@@
Substituting three intensity settings $\mu,\nu,\omega$ into Eq. 13, the lower bound of $Y_{11}^{\text{ref}}$ and the upper bound of $H_{11}^{\text{ref}}$ can be solved by Gaussian elimination, which are expressed as [39, 40]
@@ IV Key rate simulation
In the simulation, $Q_{ab}$ and $E_{ab}$ depend on Alice and Bob’s intensity settings $a$ and $b$, the total transmittance of the channel and detector $\eta$, the probability of background counts $Y_{0}$, and misalignment errors $e_{d}$ppendix B provides the calculations of $Q_{ab}$ and $E_{ab}$ using these parameters. Consistent with Refs. [35, 36], since we are considering the asymptotic regime, the key rate does not depend on the probabilities of decoy settings and basis choices. Therefore, we setting $p_{\mu}\approx 1$ and $q_{Z}\approx 1$ to maximizes key rate similar to Ref. [36]. The other simulation parameters are utilized from Ref. [42] as follows. The bidirectional error correction efficiency to $f=1.16$, the probability of background counts $Y_{0}=4\times 10^{-8}$, and misalignment errors $e_{d}=0.0108$. For the total transmittance of the channel and detector, $\eta=\eta_{d}10^{-\alpha L/10}$, where $\eta_{d}$ is the detection efficiency of the detector, $\alpha$ is the attenuation coefficient of the channel, and $L$ is the distance from Alice (Bob) to Charles. In the simulation, we set $\eta_{d}=0.53$ and $\alpha=0.2$ dB/km.
@@ V Application
It is worth noting that, according to the security proof of decoy-state protocol, the key rate obtained using the signal state, weak decoy state, and vacuum state can approximate the key rate obtained with an infinite number of decoy states [29]. Therefore, to simplify calculations while maintaining accuracy, we choose intensities of signal state $S$, weak decoy state $D_{1}$, and vacuum state $V$ in our simulation. Next, taking the signal state $S$ as an example, we analyze the relative deviation $\delta$ under intensity correlations.